I need an answer for these questions! Thanks

A small object is located 28.0 cm in front of a co
Answer #1

10) u = 28 cm

f = R/2 = 38/2 = 19 cm

v = ?

1/u + 1/v = 1/f

1/v = 1/f - 1/u

1/v = 1/19 - 1/28

v = 59.1 cm


11)

h = 42 cm

u = 1.9 m

f = -0.49 m

let v is the image distance.

apply, 1/u + 1/v = 1/f

1/v = 1/f - 1/u

1/v = 1/(-0.49) - 1/1.9

v = -0.39 m <<<<-----Answer


magnification, m = -v/u

= -(-0.39)/1.9

= 0.205